OpenStudy (anonymous):

A jewelry box has a volume of x^3 + 7x^2 + 7x – 15 cm3 and the height is x + 3 cm. Find the polynomial that would represent the area of the bottom of the jewelry box? Explain your reasoning.

OpenStudy (anonymous):

Do i do synthetic division?

OpenStudy (eyust707):

Yes

OpenStudy (anonymous):

okay after performing synthetic division, i got x^2+4x-5

OpenStudy (anonymous):

simplify it down to (x-1)(x+5)

OpenStudy (anonymous):

Would that represent the area ^^^ ?

OpenStudy (eyust707):

Yes

OpenStudy (anonymous):

How can I "explain my reasoning"?

OpenStudy (eyust707):

Well the volume of a box is: Volume = Length x Width x Height

OpenStudy (eyust707):

in this case volume = x^3 + 7x^2 + 7x – 15 and height = x + 3

OpenStudy (eyust707):

Volume/height = length x width

OpenStudy (eyust707):

so x^3 + 7x^2 + 7x – 15/x + 3 = length x width

OpenStudy (eyust707):

= the area of the bottom

OpenStudy (anonymous):

YOU ARE A GENIUS. thank yo so much @eyust707

OpenStudy (eyust707):

np =)